Consider the following sets of sample data:

A:

21,963

, 20,786, 21,202, 22,251, 21,848, 20,431, 20,336, 22,017, 20,119, 21,984, 20,261, 20,570, 20,855, 21,971

B:

86

, 74, 92, 73, 74, 83, 80, 91, 76, 99, 72

Step 1 of 2 :

For each of the above sets of sample data, calculate the coefficient of variation, CV. Round to one decimal place.
Answer #1

Solution:

a ) Given that

x x2
21963 482373369
20786 432057796
21202 449524804
22251 495107001
21848 477335104
20431 417425761
20336 413552896
22017 484748289
20119 404774161
21984 483296256
20261 410508121
20570 423124900
20855 434931025
21971 482724841
∑x=296594 ∑x2=6291484324

The sample mean is \bar x

Mean \bar x = (\sumx / n) )

=21963+20786+21202+22251+21848+20431+20336+22017+20119+21984+20261+20570+20855+21971 /14

=296594 /14

=21185.285

Mean \bar x =21185.285

The sample standard is S

  S =\sqrt{}( \sum x2 ) - (( \sum x)2 / n ) n -1

=\sqrt{}6291484324-(296594)214 /13

=\sqrt{}6291484324-6283428631.1429 /13

=\sqrt{}8055692.8571/13

=\sqrt{}619668.6813

=787.1904

he sample standard is =787.1904

Co-efficient of Variation (Sample) = Sˉ / \bar x *100%

=( 787.1904 / 21185.2857) *100%

=3.72%

Co-efficient of Variation (Sample) = 3.7%

b ) Given that

x x2
86 7396
74 5476
92 8464
73 5329
74 5476
83 6889
80 6400
91 8281
76 5776
99 9801
72 5184
∑x=900 ∑x2=74472

Solution

The sample mean is \bar x

Mean \bar x = (\sumx / n) )

=86+74+92+73+74+83+80+91+76+99+72 /11

=900 /11

=81.8182

The sample standard is S

  S =\sqrt{}( \sum x2 ) - (( \sum x)2 / n ) n -1

=\sqrt{}74472-(900)211 /10

=\sqrt{}74472-73636.3636/10

=\sqrt{}835.6364/10

=\sqrt{}83.5636

=9.1413

Co-efficient of Variation (Sample) = S / \bar x *100%

=9.141381.8182⋅100%

=11.17%

Co-efficient of Variation (Sample) =11.2%